OpenStudy (anonymous):

helpppppppppppppp please need a one on one teacher for a few more math problems

OpenStudy (anonymous):

type your problem

OpenStudy (anonymous):

sorry, that looks like a test.

OpenStudy (anonymous):

its not a test its my study guide that my teacher gave us to help us study and prepare for the test..

OpenStudy (anonymous):

its ok i understand...i was getting help but she couldnt help me finish...thanks anyway

OpenStudy (anonymous):

till need help?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

i need lots of help

OpenStudy (anonymous):

stressed out with this work

OpenStudy (anonymous):

I will try, your picture u should rotal 90degree ,help me work faster

OpenStudy (anonymous):

ok thanks let me fix it

OpenStudy (anonymous):

yea better

OpenStudy (anonymous):

thank you for your time to help me

OpenStudy (anonymous):

\[\frac{7}{2x-1}=\frac{28}{6x+14}\] here can u solve from now or u want me do it?

OpenStudy (anonymous):

can u do it please

OpenStudy (anonymous):

i been tryin to solve and answers dont look right

OpenStudy (anonymous):

\[\frac{126}{14}=x\]now u cross and mutipfly\[7\times(6x+14)=28\times(2x-1)\]\[42x+98=56x-28\]\[98+28=56x-42x\]\[126=14x\]\[9=x\]

OpenStudy (anonymous):

sorry my heybord not work properly I need rewrite it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u make this problem seem like a piece of cake

OpenStudy (anonymous):

I thimk i do Ok, the only \[\frac{126}{14}=x\]can u bring down to next the last one ,i don't know why up side down

OpenStudy (anonymous):

x=9 is right answer?

OpenStudy (anonymous):

yup you are good at this

OpenStudy (anonymous):

I start work next

OpenStudy (anonymous):

\[x^{2}-23=3x+17\]\[x^{2}-3x-23-17=0\]\[x^{2}-3x-40=0\]Factor:\[(x-8)(x+5)\] x= 8 or x =-5 the lenght can not negative than x =8 is solution

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you help me with more?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can we go another page I can get more medal?

OpenStudy (anonymous):

ok sure

OpenStudy (anonymous):

nancy i want to give u the metal

OpenStudy (anonymous):

Ab I have a jpeg for you.

OpenStudy (anonymous):

thxs